OpenStudy (anonymous):

How should i go about solving the equation: 7sin(x)+5=2cos^2(x)

jimthompson5910 (jim_thompson5910):

First start with the pythagorean identity and isolate cos^2 \[\Large \sin^2(\theta)+\cos^2(\theta) = 1\] \[\Large \cos^2(\theta) = 1-\sin^2(\theta)\]

OpenStudy (anonymous):

then distribute the 2?

jimthompson5910 (jim_thompson5910):

Yes, to get \[\Large 7\sin(x) + 5 = 2-2\sin^2(x)\]

jimthompson5910 (jim_thompson5910):

Now let \(\Large z = \sin(x)\) Squaring both sides gives you \(\Large z^2 = \sin^2(x)\)

jimthompson5910 (jim_thompson5910):

Because of those two assignments, we can replace all copies of sin(x) with z to go from \[\Large 7\sin(x) + 5 = 2-2\sin^2(x)\] to \[\Large 7z + 5 = 2-2z^2\]

OpenStudy (anonymous):

then solve normally and throw the trig back in?

jimthompson5910 (jim_thompson5910):

Exactly. You solve either by factoring, completing the square, or the quadratic formula Once you have your solutions in terms of z, you plug z = sin(x) back in and solve for x (in each equation)

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome